### v2

# UPCORE: Utility-Preserving Coreset Selection for Balanced Unlearning

Abstract
User specifications or legal frameworks often require information to be removed from pretrained models, including large language models (LLMs). This requires deleting or “forgetting” a set of data points from an already-trained model, which typically degrades its performance on other data points. Thus, a balance must be struck between removing information and keeping the model’s other abilities intact, with a failure to balance this trade-off leading to poor deletion or an unusable model. To this end, we propose UPCORE (Utility-Preserving Coreset Selection), a method-agnostic data selection framework for mitigating collateral damage during unlearning. Finding that the model damage is correlated with the variance of the model’s representations on the forget set, we selectively prune the forget set to remove outliers, thereby minimizing model degradation after unlearning. Across three standard unlearning methods, UPCORE consistently achieves a superior balance between the competing objectives of deletion efficacy and model preservation. To better evaluate this trade-off, we introduce a new metric, measuring the area-under-the-curve (AUC) across standard metrics. Our results show that UPCORE improves both standard metrics and AUC, benefiting from positive transfer between the coreset and pruned points while reducing negative transfer from the forget set to points outside of it. [^1] Correspondence to: Vaidehi Patil ¡vaidehi@cs.unc.edu¿. Code and data: https://github.com/Vaidehi99/UPCORE
1 Introduction
The widespread deployment of ML models, especially large language models (LLMs), has raised concerns around privacy, regulation, and ethical use. Trained on massive, uncurated web data, these models often memorize sensitive, copyrighted, or undesirable content (Shokri et al., 2017; Carlini et al., 2019). With regulations like the GDPR and CCPA granting individuals the “right to be forgotten,” efficient methods for removing specific data or topics from pre-trained models are increasingly necessary. Machine unlearning offers a promising solution by enabling targeted data removal without full retraining. Beyond regulatory compliance, it also helps reduce harmful outputs, protect intellectual property, and align LLMs with societal values (Jang et al., 2023). These challenges have driven renewed interest in improving model editing and unlearning techniques (Liu et al., 2024; Hase et al., 2024).
Figure 1: Left: Standard unlearning methods are applied equally to all points in the forget set. Here, outlier points in the model’s hidden space (visualized in 2D) contribute to the unintentional forgetting of points outside of the forget set (i.e. collateral damage). Right: By finding a lower-variance coreset within the forget set, UPCORE reduces damage while maintaining forget performance via positive transfer from the coreset to the pruned points.
Given the growing use of LLMs, prior work has proposed methods for removing knowledge or skills (Cao and Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022) and steering behavior in targeted ways (Sinitsin et al., 2020; Meng et al., 2022). However, such editing often induces unintended side effects, reducing utility on unrelated tasks. Effective unlearning therefore requires balancing deletion of undesired information with preservation of overall model utility. To evaluate this, current methods assess both deletion success (via “forget set” performance) and collateral effects on unrelated behaviors (via “retain set” accuracy). This is particularly important in realistic, topic-level unlearning scenarios—such as removing information about an individual or sensitive domain (Li et al., 2024)—where deletion may cause over-generalization and degrade performance on semantically or structurally similar inputs.
A key gap in existing research lies in understanding the specific data characteristics that drive over-generalization and collateral effects during unlearning. While prior work (Sheshadri et al., 2024; Chowdhury et al., 2024) has measured damage resulting from unlearning, it does not investigate how attributes of the data – such as its variance – contribute to collateral damage or whether these attributes can be controlled to optimize the trade-off between deletion efficacy and utility retention. Focusing on a topic-based setting where the forget set comprises semantically coherent groups of information, we seek to address these questions:
  1. 1.
    What measurable attributes of the forget set drive collateral effects during the unlearning process?
  2. 2.
    Can these attributes be systematically controlled to optimize the trade-off between deletion effectiveness and model utility?
We investigate which properties of the forget data correlate with collateral damage during unlearning. Our analysis reveals a strong positive correlation between the variance of the model’s hidden states corresponding to datapoints in the forget set (hidden state variance, or HSV), and the extent of collateral damage to the model after unlearning. In other words, unlearning a set of widely-distributed datapoints (as shown in Fig. 1 (left)) leads to more damage than unlearning a more densely-distributed set. Building on this insight, we hypothesize that selectively curating a coreset with lower variance from the larger forget set can help optimize this trade-off, as shown in Fig. 1 (right).
To this end, we introduce UPCORE, which constructs a core forget set by systematically identifying and pruning data points in the forget set that contribute most to the variance and thereby to collateral damage. UPCORE organizes points into an Isolation Forest (Liu et al., 2008), which identifies anomalous points in a set. By pruning these points, we reduce the variance within the forget set, which we find leads to less damage. Crucially, in addition to reducing collateral damage, UPCORE in fact leverages it by identifying two separate kinds of collateral effects: (1) Negative collateral damage: Unintended degradation of unrelated model capabilities and (2) Positive collateral transfer: The intended impact on pruned data points removed to form the core forget set. This is illustrated in Fig. 1, where pruned outlier points are still unlearned – despite not being a part of the coreset used for unlearning – due to positive transfer, and is further highlighted by our results in Table 1 and Table 3, which show that UPCORE results in better unlearning than a randomly-selected subset while also having better knowledge retention on non-forget data. We show positive collateral transfer enabled by UPCORE in Table 4, with deletion transferring to points outside the coreset. Moreover, our focus on data makes UPCORE method-agnostic: it can be applied to any data-driven unlearning framework.
We evaluate UPCORE in prompt completion and question-answering settings and across three standard unlearning methods: Gradient Ascent (Jang et al., 2023), Refusal (Ouyang et al., 2022; Maini et al., 2024) and Negative Preference Optimization (NPO) (Zhang et al., 2024b), applying each unlearning algorithm directly to the optimized core forget set obtained using UPCORE, rather than the entire forget set. We measure three critical dimensions: (1) Unlearning effectiveness, measured by the successful removal of targeted knowledge in the (a) forget set, (b) paraphrased versions of removed information as well as (c) prompts attempting to jailbreak the model.; (2) Unintended damage, where we quantify collateral effects on unrelated model capabilities; and (3) Intended transfer, where we analyze the impact on the pruned data points that were removed from the core forget set as shown Section 5.2. While we follow past unlearning work in the metrics we use to measure the trade-off between the competing objectives, we also note that the current suite of metrics measures performance at a fixed point during unlearning. This can make comparisons across methods hard, as the trade-off between deletion efficacy and model utility varies across unlearning gradient update steps. To address this, in addition to showing improvements on standard metrics, we introduce a novel set of metrics that report the area-under-the-curve (AUC) for the standard unlearning metric suite, reporting not just the performance at one fixed timestep, but measuring how a method trades off deletion with model utility across checkpoints (see Fig. 4).
Empirically, we find that across all three unlearning methods, UPCORE consistently has the highest AUC compared to baselines of unlearning on the complete forget set and choosing a random subset of forget points. In other words, UPCORE forms a Pareto frontier, maximizing unlearning effectiveness while also minimizing model damage. Moreover, UPCORE positively leverages generalization by transferring unlearning from the core set to the high-variance outlier points that were removed from the core forget set. Notably, it consistently beats baselines across all unlearning methods; this holds true across multiple metrics (e.g. ROUGE on a “retain” set, on neighborhood data closely related to (but not in) the forget set, etc.). UPCORE’s superior trade-off effectively generalizes to variations of the forgotten information, performing well on paraphrased versions of forgotten prompts as well as prompts intended to jailbreak the model. We also see these gains reflected in static evaluations of one checkpoint (as opposed to AUC, which evaluates across checkpoints); here, UPCORE obtains lower (better) ROUGE on the forget set than the random baseline while simultaneously incurring less model damage than the random and complete baselines, with the best (highest) ROUGE across all data not in the forget set.
2 Background and Related Work
Unlearning Methods for LLMs.   Machine unlearning methods fall into two categories: exact unlearning, which ensures the model is indistinguishable from one retrained without the forget data, and approximate unlearning, which modifies model parameters efficiently without full retraining. Due to the high cost of retraining LLMs, most approaches—including ours—use the latter. One class of methods trains models via RLHF to produce uninformative responses (e.g., “I don’t know”) on forget prompts (Ouyang et al., 2022; Wen et al., 2024). Yao et al. (2023) apply gradient ascent to suppress harmful outputs, substituting them with whitespace, though this causes utility loss on benign prompts. To mitigate such degradation, Chen and Yang (2023) introduce an unlearning layer effective across tasks, while Eldan and Russinovich (2023) propose a specialized architecture for removing copyrighted content. For evaluation, Maini et al. (2024) present a benchmark we adopt. Despite these advances, managing the trade-off between unlearning and utility remains challenging. We address this by introducing a data-driven coreset selection framework to minimize collateral damage.
Model Editing for Unlearning.   Model editing provides an alternative approach to unlearning by directly modifying model weights to forget target facts (De Cao et al., 2021; Dai et al., 2022; Mitchell et al., 2022; Meng et al., 2022). Following model editing work (Patil et al., 2024b), our framework employs LoRA updates with standard unlearning objectives (See Section 8.8 for more details).
Coreset Selection.   Coreset selection identifies representative subsets that preserve key dataset properties, improving computational efficiency. Given the NP-hard complexity of the exhaustive search, methods have focused on optimizing coverage, diversity, or importance (Sener and Savarese, 2018; Tan et al., 2023). By recognizing unequal contributions of data points, coreset selection has proven effective in supervised learning (Wei et al., 2015; Killamsetty et al., 2021a, b), enabling efficient performance. Our work forms new connections between these methods and the problem of unlearning in LLMs, where preserving utility and minimizing collateral damage are critical.
3 Methods
We introduce UPCORE (Utility-Preserving Coreset Design for Unlearning), an approach motivated by the observation that certain data points in the forget set disproportionately contribute to collateral damage during unlearning, primarily by increasing data variance. To address this, UPCORE reformulates pruning for core forget set selection as an outlier detection task, where outlier data points i.e. points with the greatest influence on utility degradation are identified and pruned. By minimizing variance within the forget set, UPCORE reduces unintended negative effects, ensuring more effective and targeted unlearning.
3.1 Problem Definition
Figure 2: UPCORE has four stages. First, we extract hidden states from the LLM to be modified; second, we identify outliers using Isolation Forests; third, we prune outliers to select a core forget set, and fourth, we perform unlearning on the coreset.
Let $D$ be the dataset used to train a model $M$, with $D_{F}\subset D$ representing the forget set to be unlearned. Directly unlearning $D_{F}$, i.e., applying an unlearning algorithm $\mathcal{U}$ to the model $M$ using only the data points in $D_{F}$, produces an updated model $M^{\prime}=\mathcal{U}(M,D_{F})$. However, this often leads to significant performance degradation on the retained dataset $D\setminus D_{F}$ due to over-generalization, where unlearning updates undesirably propagate beyond the forget set, impacting unrelated data points in $D\setminus D_{F}$.
To address this issue, we aim to construct a pruned forget set $D_{C}\subset D_{F}$, henceforth referred to as the core forget set, by removing points in $D_{F}$ that disproportionately drive over-generalization. The goal is to balance two competing objectives: (i) minimizing negative collateral damage, i.e., performance degradation on $D\setminus D_{F}$, and (ii) maintaining deletion accuracy, i.e., ensuring that the unlearning $D_{C}$ effectively deletes the undesirable knowledge associated with the original forget set $D_{F}$. More formally, given a damage metric $\text{Damage}_{(\mathcal{U},D_{C})}(M,M^{\prime},D\setminus D_{F})$ that quantifies the impact of unlearning $D_{C}$ on $D\setminus D_{F}$, and a deletion accuracy metric $\text{DelAcc}_{(\mathcal{U},D_{C})}(M^{\prime},D_{F})$ that evaluates the effectiveness of $\mathcal{U}$ in forgetting $D_{F}$ after unlearning on $D_{C}$, the problem can be formulated as an optimization task: $D_{C}=\arg\min_{D_{C}\subseteq D_{F}}\bigg{(}\text{Damage}_{(\mathcal{U},D_{C})}\big{(}M,M^{\prime},(D\setminus D_{F})\big{)}-\lambda\cdot\text{DelAcc}_{(\mathcal{U},D_{C})}\big{(}M^{\prime},D_{F}\big{)}\bigg{)}$ where $\lambda>0$ is a hyperparameter that controls the trade-off between the competing objectives.
3.2 Variance as a Measure of Collateral Damage
Building on prior work analyzing the cross-task generalization of forgetting methods (Zhang et al., 2024a), we investigate the relationship between attributes of the forget set $D_{F}$ and their impact on collateral damage during unlearning i.e. $\text{Damage}_{(\mathcal{U},D_{C})}\big{(}M,M^{\prime},(D\setminus D_{F})\big{)}$. Specifically, we identify the variance $Var(D_{F})$ as a critical predictor of overgeneralization. Since directly selecting optimal subsets based on unlearning performance is computationally infeasible – due to the high cost of repeatedly retraining or unlearning – our goal is to develop scalable heuristics like variance that approximate which subsets minimize collateral damage, enabling efficient and practical coreset selection for unlearning. To systematically evaluate the relationship between variance and unlearning, we analyze question-answer (QA) pairs generated from Wikipedia documents across diverse topics. Each topic-specific dataset acts as the forget set $D_{F}$, with unlearning applied separately, one topic at a time. For each forget set, we compute variance using the hidden states of the last token and the penultimate layer of the model. We then compute retain set performance as the model utility metric proposed by Maini et al. (2024), which measures performance on preserved data points after unlearning. The results, visualized in Fig. 8(a) in Section 5.4, demonstrate a strong negative correlation between HSV and model utility, indicating that variance is a potential driver of over-generalization. These findings underscore the importance of identifying and excluding points that lead to higher variance i.e. outliers to mitigate utility loss. In Section 5.4 we show similar analyses for other attributes such as model confidence and gradient similarity but find no strong correlation between utility degradation and these attributes.
3.3 UPCORE: Core Forget Set Selection
To achieve variance minimization in the forget set $D_{F}$, UPCORE frames the problem as an outlier detection task. UPCORE provides two key benefits: (1) It mitigates negative collateral damage by pruning outliers to form a more compact core forget set, and (2) It strategically exploits collateral over-generalization to extend unlearning beyond the core forget set, effectively removing the pruned points as well. As shown in Fig. 1, what might traditionally be viewed as detrimental collateral damage – when it affects points outside the forget set ($D_{F}$) – can be turned to our advantage when it impacts untrained data points within the forget set that were pruned ($D_{F}\setminus D_{C}$).
To detect these outliers, we use the Isolation Forest algorithm (Liu et al., 2008), an unsupervised learning technique that efficiently identifies anomalous data points. Isolation Forest works by recursively partitioning the dataset using random feature selections and random split values. Points that are isolated with fewer partitions, i.e. are isolated more easily, are considered outliers, as they differ significantly from the majority of the data. This makes the Isolation Forest algorithm particularly effective for high-dimensional data where traditional distance-based outlier detection methods may fail. These outliers, characterized by their isolation in the feature space, are likely to contribute to high variance and over-generalization during the unlearning process. UPCORE proceeds as follows (illustrated in Fig. 2):
Stage 1: Hidden Feature Extraction: We extract hidden state representations $\mathcal{H}$ from the model’s penultimate layer, as it typically encodes high-level semantic abstractions while retaining generalization capacity, unlike the final layer which is often biased toward task-specific outputs (Skean et al., 2025) (See Fig. 2 left), corresponding to the final token of each question in $D_{F}$. These representations, which reflect the model’s internal representation of the data, serve as input features for outlier detection. This step is guided by our analysis in Section 3.2, which highlights the strong link between hidden state variance and collateral damage.
Stage 2: Training the Isolation Forest and Computing Anomaly Scores: We train an Isolation Forest model $\mathcal{I}$ on the forget set $D_{F}$ to model its distribution, recursively partitioning the data to detect outliers (see Fig. 2 middle). Points isolated more quickly and requiring fewer splits are flagged as outliers, indicating disproportionate contributions to variance in the hidden state space. For each $d\in D_{F}$, $\mathcal{I}$ assigns an anomaly score $\text{score}(d)$ based on the average path length $h(d)$ required to isolate the point across an ensemble of binary trees. Shorter path lengths correspond to higher anomaly scores, indicating points that contribute to variance and thus collateral effects. Additional details are provided in Section 7.
Stage 3: Prune Outliers and Setting Stopping Criterion: To construct the pruned coreset $D_{C}$, we apply a threshold $\tau$ on the anomaly scores from the Isolation Forest model: $D_{C}=\{d\in D_{F}\mid\text{score}(d)\leq\tau\}$. Data points with scores above $\tau$ are excluded as outliers, as they disproportionately contribute to variance. We hypothesize that removing these outliers will reduce utility degradation while preserving core information for forgetting. The threshold $\tau$ is determined via a stopping criterion, which can be chosen as follows: (1) Coreset Size Control: Specify a desired coreset size $|D_{C}|$ to ensure an appropriate number of inliers (2) Proportional Pruning: Select the top $k\%$ of points with the lowest anomaly scores to maintain a consistent pruning ratio. By selecting $\tau$ based on user requirements, UPCORE provides fine-grained control over the trade-off. In practice, we prune 10% of the data points in our main experiments and additionally conduct scaling experiments that vary the pruning percentage (see Section 8.1) to analyze its impact on the trade-off dynamics.
Stage 4: Unlearning on the Coreset: After selecting the pruned coreset $D_{C}$, UPCORE applies the unlearning algorithm $\mathcal{U}$ to the model $M$, resulting in $M^{\prime}_{\textsc{UPCORE}{}}=\mathcal{U}(M,D_{C})$ (See Fig. 2 end). This process removes the influence of $D_{F}$ while minimizing utility degradation on $D\setminus D_{F}$. By focusing on $D_{C}$, UPCORE ensures targeted unlearning and positively leverages collateral effects, as unlearning $D_{C}$ also deletes much of $D_{F}$’s influence, even those parts not explicitly included in $D_{C}$. We later show this positive transfer empirically in Table 4.
4 Experimental Setup
Unlearning Methods and Baselines.   We test UPCORE with three standard unlearning methods: gradient ascent, refusal, and negative preference optimization, applied to a Llama-3.1-8B (Dubey et al., 2024) base model. In all cases, models are trained using both the forget set (complete or sampled) and a retain set, which contains examples of data that should not be forgotten, providing a contrastive signal. We consider the following unlearning methods:
  • Gradient Ascent (Jang et al., 2023): Gradient Ascent maximizes the training loss on the forget set $D_{F}$. For each $x\in D_{F}$, the objective is to maximize the loss.
  • Refusal (Ouyang et al., 2022): Refusal trains the model to respond to sensitive prompts with neutral, non-informative answers, such as “I don’t know.
  • Negative Preference Optimization (NPO) (Zhang et al., 2024b): NPO is a stable form of DPO (Rafailov et al., 2024) designed for unlearning. It reduces the gap between the likelihood of the target data and the likelihood from the original model while ensuring the unlearned model remains closely aligned with the original (See Section 8.7 for more details).
We evaluate UPCORE, which is a dataset selection method, against three other selection methods as baselines: (1) unlearning applied to the entire forget set (i.e. no selection), and (2) unlearning performed on a randomly subsampled subset of the forget set, matched in size to the coreset curated by UPCORE (i.e. random selection) (3) $D^{2}$-pruning (Maharana et al., 2024), a standard coreset selection method that selects the coreset based on example diversity and difficulty.
We evaluate on factual questions across two settings: prompt completion and question answering, described below, and we include further details on these settings in Section 8.5.
4.1 Metrics and Answer Extraction
Following prior work (Maini et al., 2024), we evaluate models using a suite of metrics. We compute ROUGE (Lin, 2004) between reference and model answers to assess both utility and deletion effectiveness, as ROUGE captures content overlap in factual QA, unlike classification-based metrics with fixed labels. To measure unintended model damage, we report ROUGE on the retain set, neighborhood data, and Real World / Real Authors datasets (Maini et al., 2024), where higher is better. For deletion, we compute ROUGE on the forget set (lower is better) and on pruned forget examples to assess positive collateral transfer. We also report model utility, defined as the harmonic mean of the normalized conditional probabilities $P(a\mid q)^{1/|a|}$, following Cho et al. (2014), and the truth ratio, which compares the likelihood of correct vs. incorrect answers (Maini et al., 2024). We also evaluate all metrics on paraphrased and jailbreak variants of the forget set, where paraphrased variants reword target examples (Krishna et al., 2023) and jailbreak variants probe adversarial prompts designed to bypass unlearning (Zou et al., 2023; Jin et al., 2024).
AUC Metric.
While ROUGE and model utility are standard unlearning metrics, they provide only a single-point snapshot of model performance. This is limiting, as unlearning inherently involves a tradeoff between forgetting and model damage that evolves over training steps. [^2] We use steps throughout; in main results, 1 epoch = 50 steps. Early checkpoints may retain higher utility but underperform on deletion, while later ones improve forgetting at the cost of increased damage (Fig. 4). Since the number of unlearning steps varies across works, direct comparisons become difficult. To address this, we propose evaluating unlearning as a dynamic tradeoff over time. Instead of reporting ROUGE Forget and ROUGE Retain at one step, we compute the area under the curve (AUC) between these metrics over unlearning steps. This tradeoff is visualized in Fig. 4, which plots inverse ROUGE on forget data (X-axis) versus ROUGE on neighboring data (Y-axis). [^3] This curve differs from Table 1: we compute AUC before averaging across topics here, and after in Table 1. We construct Pareto curves comparing deletion metrics (e.g., ROUGE Forget) with utility metrics (e.g., ROUGE Retain, ROUGE Neighborhood) across steps, and use the AUC as a unified, global measure of performance. We also study the effect of evaluation granularity on AUC (Table 10), and show that AUC correlates with overall unlearning effectiveness (Table 3) and negatively with forget data variance (Table 7). In Section 5.4 we analyze the effect of varying the number of steps on this AUC metric and find our results are stable across different granularities. Therefore we adopt a standard of 50 steps, corresponding to one epoch.
5 Experimental Results and Discussion
5.1 UPCORE Balances Deletion and Model Utility
Design.   To assess whether UPCORE improves the Pareto frontier between deletion effectiveness and utility retention, we compare AUC values of models trained with UPCORE and baseline methods (Section 4). AUC is computed using: (1) Deletion Effectiveness, measured as $(1-\text{ROUGE})$ on the forget set (X-axis), and (2) Utility Retention, measured via ROUGE on non-forget datasets, including neighborhood data and the aggregated model utility metric from Maini et al. (2024) (Y-axis). We perform evaluations are performed on both Counterfact and TriviaQA topics.
Results.   Fig. 4 illustrates the AUC metric, which quantifies the trade-off between forget performance and model utility by measuring the area under the Pareto frontier. A higher AUC reflects better balance – more effective forgetting with less utility loss. UPCORE slows utility degradation by unlearning on a variance-based coreset. Table 1 shows that across the Counterfact dataset and three unlearning methods, UPCORE consistently outperforms baselines – complete forget set and random subsample – by 3 to 7 AUC points, confirming its method-agnostic effectiveness. Table 2 extends this to TriviaQA (using gradient ascent) (see Table 6 for results using other unlearning methods on TriviaQA), where UPCORE maintains its advantage with up to 3 AUC points improvement. Finally, Table 3 reports ROUGE and utility scores at epoch 10, showing that UPCORE achieves the highest utility and non-forget ROUGE, while preserving competitive forget performance. We report p-values in Appendix Table 12, finding the gains are statistically significant.
Table 1: AUC across the two competing objectives: (1) Deletion Effectiveness, defined as $(1-\text{ROUGE})$ on the forget set (X-axis), and (2) Model Utility, averaged across Counterfact topics and evaluated via ROUGE scores on multiple utility datasets, including neighborhood data and an aggregate model utility across datasets (Y-axis). We compare three unlearning methods: Gradient Ascent, Refusal, and NPO. Error bars indicate standard deviation across 3 seeds.
Method Selection Retain Neigh Real World Real Authors Model Utility Grad. Ascent Complete 0.488 $\pm$ 0.015 0.568 $\pm$ 0.018 0.720 $\pm$ 0.016 0.891 $\pm$ 0.020 0.343 $\pm$ 0.012 Random 0.495 $\pm$ 0.017 0.558 $\pm$ 0.016 0.731 $\pm$ 0.015 0.907 $\pm$ 0.019 0.353 $\pm$ 0.014 $D^{2}$-pruning 0.493 $\pm$ 0.016 0.552 $\pm$ 0.017 0.723 $\pm$ 0.016 0.920 $\pm$ 0.018 0.349 $\pm$ 0.013 UPCORE 0.523 $\pm$ 0.008 0.608 $\pm$ 0.010 0.769 $\pm$ 0.009 0.933 $\pm$ 0.011 0.387 $\pm$ 0.007 Refusal Complete 0.493 $\pm$ 0.016 0.488 $\pm$ 0.017 0.714 $\pm$ 0.015 0.890 $\pm$ 0.018 0.366 $\pm$ 0.014 Random 0.456 $\pm$ 0.015 0.458 $\pm$ 0.016 0.644 $\pm$ 0.014 0.819 $\pm$ 0.017 0.332 $\pm$ 0.013 $D^{2}$-pruning 0.473 $\pm$ 0.013 0.478 $\pm$ 0.014 0.632 $\pm$ 0.011 0.805 $\pm$ 0.015 0.341 $\pm$ 0.010 UPCORE 0.500 $\pm$ 0.007 0.524 $\pm$ 0.009 0.744 $\pm$ 0.008 0.920 $\pm$ 0.010 0.381 $\pm$ 0.006 NPO Complete 0.281 $\pm$ 0.014 0.237 $\pm$ 0.015 0.192 $\pm$ 0.013 0.342 $\pm$ 0.017 0.199 $\pm$ 0.012 Random 0.253 $\pm$ 0.015 0.271 $\pm$ 0.014 0.195 $\pm$ 0.013 0.308 $\pm$ 0.016 0.186 $\pm$ 0.011 $D^{2}$-pruning 0.265 $\pm$ 0.013 0.254 $\pm$ 0.014 0.193 $\pm$ 0.012 0.320 $\pm$ 0.016 0.190 $\pm$ 0.011 UPCORE 0.329 $\pm$ 0.006 0.319 $\pm$ 0.008 0.246 $\pm$ 0.007 0.414 $\pm$ 0.009 0.248 $\pm$ 0.005
Figure 3: Trading-off between deletion effectiveness and model utility forms a Pareto frontier across steps, shown here averaged across Counterfact topics with Gradient Ascent.
Figure 4: AUC between forget set ROUGE and neighborhood data ROUGE averaged across topics in Counterfact. UPCORE reduces damage to neighborhood data.
Table 2: Evaluation metrics from Table 1 shown for Gradient Ascent on the TriviaQA topics. Error bars indicate standard deviation across 3 seeds.
Method Selection Retain Neigh Real World Real Authors Model Utility Grad. Ascent Complete 0.153 $\pm$ 0.004 0.285 $\pm$ 0.005 0.226 $\pm$ 0.004 0.155 $\pm$ 0.003 0.135 $\pm$ 0.004 Random 0.159 $\pm$ 0.005 0.304 $\pm$ 0.006 0.222 $\pm$ 0.005 0.157 $\pm$ 0.004 0.136 $\pm$ 0.004 $D^{2}$-pruning 0.162 $\pm$ 0.003 0.310 $\pm$ 0.004 0.224 $\pm$ 0.003 0.157 $\pm$ 0.003 0.141 $\pm$ 0.003 UPCORE 0.165 $\pm$ 0.002 0.318 $\pm$ 0.003 0.227 $\pm$ 0.002 0.158 $\pm$ 0.002 0.147 $\pm$ 0.002
Table 3: ROUGE scores and model utility across topics from the Counterfact dataset for a fixed epoch of Gradient Ascent. UPCORE consistently has higher performance on data outside the forget set, with the least degradation among methods and closest performance to the base model, while still having a high forget rate.
Method Forget Retain Neigh. Real Authors Real World Model Utility
Base model 0.997 0.546 0.820 1.000 0.872 0.433
Complete 0.018 0.381 0.144 0.669 0.446 0.182
Random 0.011 0.411 0.104 0.724 0.499 0.211
UPCORE 0.017 0.430 0.190 0.706 0.528 0.350
5.2 Positive and Negative Transfer
Design.   Here, we measure both positive and negative transfer. To assess whether unlearning on the core forget set induces deletion in the pruned data points (positive collateral transfer), we measure the ROUGE score of the unlearned model on these points. A significant drop in ROUGE would indicate that the forgetting process extends beyond the explicitly unlearned subset. We measure negative transfer on the neighborhood data, examining the AUC between ROUGE on the neighborhood datapoints and forget set ROUGE.
Table 4: ROUGE score on pruned datapoints. Both for UPCORE and random sampling, unlearning on a subset of datapoints translates to other datapoints not in the subset.
Method Random UPCORE
Base model 1.000 1.000
Gradient Ascent 0.022 0.053
Refusal 0.169 0.127
NPO 0.206 0.231
Results.   As shown in Table 4, ROUGE on pruned points drops from 1.00 to 0.053 (Gradient Ascent) and 0.127 (Refusal), indicating that unlearning transfers to pruned points despite not being directly targeted—likely due to topic-level over-generalization. This transfer is not unique to UPCORE; a similar drop occurs with a random subsample of the same size, suggesting that pruned points share a semantic neighborhood with the forget set and are thus indirectly affected. However, in terms of negative transfer, UPCORE achieves substantially higher AUC on neighborhood data (Fig. 4), indicating reduced unintended damage compared to random sampling. These findings are further supported by utility AUC gains in Table 1, highlighting that while unlearning generalizes well across topics in the positive direction, variance-based pruning better limits collateral damage in the negative direction. We hypothesize that the gains on Counterfact are higher than those on TriviaQA (see Table 1 and Table 2) due to the latter’s higher semantic density, which facilitates stronger positive transfer across related examples.
5.3 Robustness to Jailbreaks
Design.   To evaluate robustness against blackbox attacks, we test whether unlearning on the core forget set generalizes to adversarial/jailbreak prompts designed to elicit the same information (see Section 8.6 for examples and generation details). We report the AUC of (1-ROUGE) and ROUGE on non-forget data (e.g., retain set, neighborhood) in Table 5. Higher AUC indicates greater robustness against extraction attacks (Zou et al., 2023; Jin et al., 2024). We repeat this analysis with paraphrased variants in Appendix Table 13, with similar findings in terms of robustness to paraphrased queries.
Table 5: Evaluation metrics from Table 1 averaged across topics in Counterfact shown for Gradient Ascent, assessed for robustness to jailbreak variants of the forget data with the same utility data.
Method Selection Retain Neigh Real World Real Authors Model Utility
Jailbreak Complete 0.417 0.474 0.599 0.743 0.291
Random 0.430 0.470 0.629 0.787 0.305
UPCORE 0.455 0.512 0.665 0.819 0.335
Results.   As shown in Table 5 and Table 13, UPCORE achieves higher AUC across settings and utility datasets, outperforming baselines even under rephrases and jailbreak attacks. This indicates a superior trade-off and suggests that positive transfer from the core forget set generalizes to input variations eliciting the same target information.
5.4 Correlation of Hidden-State Variance and Utility.
Figure 5: Model utility and hidden state variance (forget data) show a strong negative correlation (-0.714) across data from multiple topics.
We show that hidden state variance is strongly negatively correlated with model utility after unlearning (Pearson correlation = -0.714), supporting its role as a reliable proxy for unlearning-induced degradation. Additionally, Section 8.9.1 demonstrates a consistent inverse relationship between variance and AUC across topics, reinforcing that minimizing variance leads to improved deletion-utility trade-offs. As illustrated in Fig. 5 (top left), we plot the hidden state variance of different data clusters against the utility metric from Maini et al. (2024) after a fixed number of unlearning steps, and observe a clear negative correlation.
5.5 Additional Results Summary
We briefly describe a number of additional results and analyses included in the appendix.
two are strongly negatively correlated with a Pearson correlation of -0.714.
Alternate Outlier Detection Methods.   In Section 8.9.2, we empirically compare multiple outlier detection methods and find that Isolation Forest achieves the highest AUC, indicating its superior ability to identify informative outliers for pruning.
Effect of Forget Set Size.   Section 8.9.3 shows that UPCORE retains its effectiveness even when the forget set is reduced to 50% of its original size, maintaining improved utility relative to baselines; however, we hypothesize a lower bound below which pruning may become ineffective due to limited signal.
Granularity of AUC Steps.   Section 8.9.5 examines the impact of varying the number of unlearning steps used to compute AUC and finds the results to be stable across different granularities. This motivates our choice of epochs for the main results in Section 5.
Coreset Size.   In Section 8.1 we examine the effect of scaling the size of the coreset, increasing the number of outliers pruned. We find that different pruning percentages generally lead to similar AUCs, as higher pruning leads to less model damage but also less forgetting.
6 Conclusion
We propose UPCORE, a utility-preserving coreset selection framework for unlearning in LLMs that minimizes collateral damage while ensuring effective deletion. Empirically, we find that hidden state variance in the forget data strongly influences utility degradation. By pruning high-variance outliers to form a core forget set, UPCORE improves the trade-off between deletion and retention. We quantify this trade-off using area-under-the-curve across unlearning steps. Results show that UPCORE substantially reduces unintended performance loss and can be combined with any data-driven unlearning method, offering a principled and generalizable approach to utility-aware unlearning.
Acknowledgments
We would like to thank Peter Hase for his feedback on an initial draft of the paper. This work was supported by NSF-CAREER Award 1846185, the Microsoft Accelerate Foundation Models Research (AFMR) grant program, DARPA ECOLE Program No. HR00112390060, and NSF-AI Engage Institute DRL-2112635. Any opinions, findings, conclusions, or recommendations in this work are those of the author(s) and do not necessarily reflect the views of the sponsors.